Jennaofcbaby:

Math question @Nish15

Sailor:

So where's the question?

Jennaofcbaby:

15x-5 and that’s negativity

Jennaofcbaby:

@Nish15

Sailor:

Does 15x-5 equal anything?

Jennaofcbaby:

=7

CripQUEZZ:

oh ok

NFT0X:

well lets see, whats 15 x 5?

Sailor:

Ok. So first, you would add 5 to 7

Jennaofcbaby:

@sailor wrote:
Ok. So first, you would add 5 to 7
Okk

Sailor:

Then divide that by 15 and you'll have the answer.

Jennaofcbaby:

@sailor wrote:
Then divide that by 15 and you'll have the answer.
Oh got it thankss

Sailor:

@jennaofcbaby wrote:
@sailor wrote:
Then divide that by 15 and you'll have the answer.
Oh got it thankss
No problem

CripQUEZZ:

15x-5=7 first add 5 to 7 to cancel out the -5 which equalys 15x = 12 now divide 15 from both sides because u want ur variable by its self so 15x/15 = 12/15 so now x is by itself and it equals x = 12/15